OpenStudy (winner.mo7):

What is the equation of the line with y-intercept=3 and slope =2/3?

OpenStudy (anonymous):

y intercept =3 means point is (0,3) now as equation of line through a point x1,y1 and slope m is given by y-y1=m(x-x1) so the equation of line through (0,3) and slope 2/3 is (y-3)=2/3(x/0) i hope u can simplify it further

OpenStudy (anonymous):

the equation will be y-3=2/3(x-0) not x/0 sorry for mistake

OpenStudy (haseeb96):

y = (slope)(x) + (y intercept) Since this is the formula of an line, you can simply put up values to get it. y = (2/3)(x) + 3

OpenStudy (winner.mo7):

Thank you very much :)